The t statistic for a test of ?0:?=8 H 0 : μ = 8 ??:?>8 H A : μ > 8 based on n = 17 observations has the value t = 1.1. (a) What are the degrees of freedom for this statistic? (b) Using the appropriate table in your formula packet, bound the p-value as closely as possible: < p-value
Answer #1

H_0:\mu=8

H_a:\mu>8

a) degree of freedom = n-1=17-1=16

b) This is a right tailed test

test statistic t = 1.1

degree of freedom = 16

P value for the test = 0.1438 ( using excel formula =T.DIST.RT(1.1,16) )

0.1<p-value